# Date of publication xxxx 00, 0000, date of current version xxxx 00, 0000. xx

### v2

@@ VI-B3 Processing latency
VI-B4 Visual Demonstration of the Proposed Approach
@@
VII Future Work and Limitations
This research identified two main limitations. First, when the robot approaches the object within approximately 60 cm, the object is no longer detected. This distance threshold was determined experimentally. At distances below 60 cm, the object’s reflection peak becomes too close to the direct-path signal, causing it to be overshadowed. Additionally, multipath components generated by the direct signal may further complicate the separation of the object’s peak. While this effect is primarily due to antenna design and the dominance of the direct path, the object’s material properties also influence the outcome. Highly reflective materials generate stronger peaks, whereas absorptive materials produce weaker responses that are more easily masked. To ensure reliable detection at closer distances, improved antenna isolation between the transmitter and receiver could help reduce the direct-path signal, or higher bandwidth settings could be used to achieve better timing resolution. Increasing the attenuation between the TX antenna and RX antenna can also help reduce direct-path interference, for example by utilizing dedicated UWB antenna arrays. Second, the analysis showed that peak prominence decreases and width increases with distance, however, the filters used in this work employed fixed thresholds. In future work, adaptive thresholding methods could be implemented to account for distance-dependent variations in peak characteristics, thereby improving detection robustness across different ranges and reducing false positives and false negatives.
@@ VIII Conclusion
This paper presented an infrastructure-free UWB radar-based approach for obstacle mapping on mobile robots. The proposed method consists of three steps: (i) target identification (based on CIR peak detection), (ii) filtering (based on peak properties, signal-to-noise score, and phase-difference of arrival), and (iii) clustering (based on distance estimation and angle-of-arrival estimation). The experimental evaluation across channels 5 and 9 and three material types demonstrated that the proposed system achieved high distance-estimation accuracy and detection accuracy, with obstacle localisation errors below 11 cm and precision above $90\%$ on channel 9. The filtering methods and clustering effectively suppressed noise and multipath reflections without requiring any prior background data, making it well suited for dynamic environments. By integrating AoA estimation, the approach enables object localisation and obstacle mapping, providing a strong foundation for use in SLAM and other real-time robotic applications.
[Uncaptioned image] Adelina Giurea received the M.Sc. degree in Information Engineering Technology from Ghent University, Belgium, in September 2025. She is currently pursuing the Ph.D. degree with the IDLab Research Group, focusing on UWB and BLE.
[Uncaptioned image] Ir. Stijn Luchie was born in Veurne in 1999. He received his M.Sc. degree in Computer Science Engineering from Ghent University, Belgium, in September 2022. Shortly thereafter, in October 2022, he started as a researcher with the Department of Information Technology (INTEC) at Ghent University in the IDLab research group. His primary area of scientific interest centres on ultra-wideband technology, with a specific focus on the self-calibration of anchor networks and the applications of UWB radar.
[Uncaptioned image] Dr. ir. Dieter Coppens received the master’s degree in electrical engineering from Ghent University, Belgium, in 2021 and the Ph.D. degree with the IDLab Research Group in 2025. His research interests are wireless networking, indoor localization system based on Ultra-wideband technology and machine learning.
[Uncaptioned image] Prof. dr. ir. Jeroen Hoebeke received the Masters degree in Engineering Computer Science from Ghent University in 2002. In 2007, he obtained a Ph.D. in Engineering Computer Science with his research on adaptive ad hoc routing and Virtual Private Ad Hoc Networks. Current, he is an associate professor in the Internet Technology and Data Science Lab of Ghent University and imec. He is conducting and coordinating research on wireless (IoT) connectivity, embedded communication stacks, deterministic wireless communication and wireless network management. He is author or co-author of more than 200 publications in international journals or conference proceedings.
[Uncaptioned image] Prof. dr. ir. Eli De Poorter is currently a Professor with the IDLab Research Group, Ghent University, and imec. His team performs research on wireless communication technologies such as (indoor) localization solutions, wireless IoT solutions, and machine learning for wireless systems. He performs both fundamental and applied research. For his fundamental research, he is currently the coordinator of several research projects (SBO, FWO, and GOA). He has over 200 publications in international journals or in the proceedings of international conferences. For his applied research, he collaborates with industry partners to transfer research results to industrial applications, and to solve challenging industrial research problems.
[…]

### v3

@@ VI-B4 Processing latency
VI-B5 Visual Demonstration of the Proposed Approach
@@
VII Discussion
This research identified two main limitations. First, objects closer than approximately 60 cm are not detected. In this near-field region, the reflection peak merges with the strong direct path between transmitter and receiver, making separation impossible. This effect is primarily caused by insufficient isolation between the antennas, but material properties further influence detectability. Highly reflective objects still generate visible peaks, while absorptive materials are easily masked. Improved antenna isolation between the transmitter and receiver could help reduce the direct-path signal, or higher bandwidth settings could be used to achieve better timing resolution. Although this blind zone prevents near-field detection, objects are consistently detected at larger distances, allowing obstacle avoidance or SLAM systems to react before entering the 60 cm range. For tasks requiring the robot to operate closer, this limitation can be handled by fusing UWB radar data with a close-range sensor. Second, the analysis showed that peak prominence decreases and width increases with distance, however, the filters used in this work employed fixed thresholds.
@@ VIII Conclusion
This paper presented an infrastructure-free UWB radar-based approach for obstacle mapping on mobile robots. The proposed method consists of three steps: (i) target identification (based on CIR peak detection), (ii) filtering (based on peak properties, SNR-based score, and phase-difference of arrival), and (iii) clustering (based on distance estimation and angle-of-arrival estimation). The experimental evaluation across channels 5 and 9 and three material types demonstrated that the proposed system achieved high distance-estimation accuracy and detection accuracy, with obstacle localisation errors below 17 cm. Even in challenging low-reflectivity scenarios such as concrete, the method achieves a precision of 73.42% and a recall of 83.38% on channel 9. The approach enables lightweight, anchor-free mapping and provides a solid foundation for future UWB-based SLAM and other robotic applications. Remaining challenges include detection at close range (¡ 60 cm) and improving antenna isolation to reduce direct-path interference. Future work may explore adaptive thresholds for peak properties, improved antenna designs, or fusion with complementary short-range sensing to address these limitations.